### v6

# A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:

A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
IMPLICATIONS FOR TAX COMPETITION***Andreas Haufler, Seminar for Economic Policy, LMU Munich, Akademiestr. 1, 80799 Munich, Germany. Email: Andreas.Haufler@econ.lmu.de.  Hayato Kato, Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan. Email: hayato.kato@econ.osaka-u.ac.jp.
 Financial support from the German Research Foundation through CRC TRR 190 (Haufler) and the JSPS KEKENHI (JP20H01495, JP22K13388, JP24H00014) (Kato) is gratefully acknowledged. This study is conducted as a part of the Project “Economic Policy Issues in the Global Economy” undertaken at the Research Institute of Economy, Trade and Industry (RIETI). Kato thanks the University of Munich for their hospitality while part of this paper was written. We thank Tomohiro Ara, Real Arai, Ruby Doeleman, Antoine Ferey, Emanuel Hansen, Jim Hines, Niels Johannesen, Makoto Hasegawa, Jota Ishikawa, Mick Keen, Sang-Hyun Kim, Yoshimasa Komoriya, Jonas Löbbing, Mohammed Mardan, Jakob Miethe, Kenichi Nishikata, Hirofumi Okoshi, Pascalis Raimondos, Dirk Schindler, Karl Schulz, Georg Thunecke, and Alfons Weichenrieder for many helpful comments and discussions. Thanks also to seminar and conference participants at the IIPF in Utah State, MPI Berlin, Goettingen, Osaka, Kyoto, Shimane, QUT Brisbane, Ryukoku, Gakushuin, CESifo (Munich & Venice Summer Institute), and RIETI for helpful comments.
Andreas Haufler and Hayato Kato
The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms. We analyse tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT. Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals. However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals. Our calibration exercise shows that introducing a GMT rate of 15% results in a regime where only the havens adopt split tax rates. Welfare and tax revenues fall in the havens but rise in the non-haven, yielding a positive net gain worldwide.
Profit shifting by multinational enterprises (MNEs) remains a central challenge for international corporate taxation. Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., , 2023), a pattern also documented in microdata for France (Davies et al., , 2018), the UK (Bilicka, , 2019), and the US (Guvenen et al., , 2022) In response to the resulting revenue losses, the OECD launched its Base Erosion and Profit Shifting (BEPS) initiative (OECD, , 2013), whose core outcome is the introduction of a global minimum tax (GMT), also known as Pillar Two. [^1] See Fuest and Neumeier, (2023); Hebous et al., (2024) for the institutional background. Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., , 2025; Ogawa and Tsuchiya, , 2025; Richter, , 2025). In 2021, more than 130 countries including many tax havens agreed on a 15% GMT; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT. [^2] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed October 1, 2025). Despite announcing after the June 2025 G7 summit that it would not fully comply with the GMT, the US had already implemented a similar tax on Global Intangible Low-Tax Income under the 2017 Tax Cuts and Jobs Act (Chodorow-Reich et al., , 2024).
A limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years. [^3] Since 2016, such groups have been required to file country-by-country reports on their activities, profits, and taxes in each country of operation (Doeleman et al., , 2024, Section 2). The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, , 2022, p.14). According to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT. [^4] See Supplemental Appendix A. This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, , 2020, # 505, p. 233). Nevertheless, over 300 billion EUR of MNEs’ profits remain outside its scope. Moreover, Orbis tends to oversample large firms (Bajgar et al., , 2020), so the 90% coverage rate should be viewed as an upper bound.
Does this incomplete coverage affect the effectiveness of the GMT? It does not if low-tax countries apply the GMT rate to all firms regardless of size. It does matter, however, if they differentiate their tax rates, applying the GMT rate to firms above the threshold and a lower rate to those below. Such a split is inherent in the regulation: tax havens may implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a top-up tax. [^5] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT) (Devereux, , 2023). This system has been adopted in Bahrain, Ireland, Switzerland and the UAE, among other countries with base statutory corporate tax rates below 15% (PwC, , 2025).
Against this background, we ask when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT. Does a 15% GMT rate induce such a split, and how large are the resulting revenue and welfare gains? How do these outcomes change with broader GMT coverage?
To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits to tax-haven countries. Countries maximise welfare—a weighted sum of tax revenues and, for the non-haven, the post-tax income of MNEs’ owners—and compete for the MNEs’ profits. In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rates for all MNEs. In the second stage, countries set tax rates, either uniform or split, with large MNEs taxed at least at the GMT rate. The final stage involves MNEs’ profit-shifting decisions.
Our analysis shows that introducing the GMT with partial coverage generates a sequence of non-cooperative tax equilibrium regimes between the non-haven and the havens. When the GMT rate is binding but low, the havens commit to it for all MNEs to raise the non-haven’s uniform tax rate. As the GMT rate rises, the havens begin to split their rates, applying the GMT to large MNEs and a lower rate to small MNEs. For sufficiently high GMT rates, the non-haven also splits its rate, as limited competition for large MNEs’ profits shifts the focus towards small MNEs.
We calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms. A 15% GMT rate induces the havens to split their tax rates, which is consistent with observed responses in some tax havens including Ireland and Switzerland. It benefits the non-haven while harms the havens, generating net global revenue (welfare) gains of 136 billion USD (83 bUSD), comparable to existing estimates (Baraké et al., , 2022; Hugger et al., , 2024). If the GMT covers all firms, the 15% rate does not produce a split tax system but increases revenues and welfare in the non-haven and globally even more. This global revenue gain would have to be set against the additional compliance costs for multinationals, however.
We contribute to the recent theoretical studies on tax competition under a GMT. They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, , 2022; Hebous and Keen, , 2023); real economic activities (Janeba and Schjelderup, , 2023; Schjelderup and Stähler, , 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, , 2025). All these models assume the GMT applies uniformly, ignoring the firm-size threshold below which firms are exempt from the GMT. Our model highlights this threshold, allowing countries to tax firms above and below it differently. [^6] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, , 1999; Keen, , 2001; Bucovetsky and Haufler, , 2007; Gaigné and Wooton, , 2011).
Our simple quantitative exercise complements recent evaluations of corporate taxation using trade and multinational models (Wang, , 2020; Ferrari et al., , 2023; Shen, , 2024; Dyrda et al., , 2024), where taxes are assumed to be exogenous and/or uniform across firms. By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes.
Two empirical papers (Bachas et al., , 2025; Clifford et al., , 2025) closely related to ours examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively. Their findings that GMT revenue gains depend strongly on its coverage emphasizes the relevance of our analysis.
1 The model
1.1 Setup
We consider tax competition between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. There is a continuum of heterogeneous MNEs which are headquartered in the non-haven and have affiliates in all havens. Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven. [^7] Firm heterogeneity could result from (executives’) productivity (e.g., Melitz, , 2003; Lippi and Schivardi, , 2014). MNEs can shift some of these profits to the havens to maximise their post-tax profits. In the following, we assume collective decision-making by the havens. This assumption is relaxed in the quantitative analysis later (see also footnote 16) and further examined in Supplemental Appendix B.
We assume the non-haven and all havens have signed the GMT agreement and must set tax rates no lower than the GMT rate $t_{M}$. [^8] Devereux, (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates. The GMT applies only to MNEs with profits $\pi\geq\pi_{M}$ (‘large MNEs’), allowing countries to set lower rates for firms with $\pi<\pi_{M}$ (‘small MNEs’).
The non-haven maximises a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of MNEs, which are all owned by residents of the non-haven. Since havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
We analyse a three-stage game. In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs—at or above the GMT rate—or to allow differentiated rates. [^9] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 6 and in related work on endogenous timing in games (Huck et al., , 2002; Kempf and Rota-Graziosi, , 2010). The choice between a uniform and a split tax system is more fundamental, and less frequently revised, than the choice of tax rates. The havens decide collectively on whether to unify or split their rates. In the second stage, countries set their tax rates non-cooperatively. Those that committed choose a single rate; others set different rates for large and small MNEs. In the final stage, MNEs shift profits given these tax rates. The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium, referred to as the tax-competition equilibrium.
1.2 Profit shifting by multinationals
Each MNE shifts a share $\theta_{h}$ of its pre-tax profits $\pi$ from its non-haven headquarters to each tax haven affiliate $h$, and incurs transaction costs $C_{h}(\cdot)$ for doing so. These costs can be interpreted as concealment costs or as expected fines, if detected. We assume the costs are quadratic in the level of profit shifting but decrease with the size of a firm: $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$. [^10] This specification is standard in the literature (e.g., Hines and Rice, , 1994; Huizinga and Laeven, , 2008; Suárez Serrato, , 2018; Bilicka et al., , 2024). The parameter $\delta$ captures the difficulty of profit shifting. [^11] This can be higher, for example, due to taxpayer audit (Lediga et al., , 2023).
Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^12] The GMT tax base equals that of national corporate taxes, ignoring additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler, (2024).
$\max_{\{\theta_{h}\}_{h=1}^{H}}\ \widetilde{\pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)=(1-t_{n})\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi+\sum_{h=1}^{H}(1-t_{h})\theta_{h}\pi-\sum_{h=1}^{H}\frac{\delta(\theta_{h}\pi)^{2}}{2\pi}.$ (1)
The optimal choice is
$\theta_{h}=\frac{t_{n}-t_{h}}{\delta},$ (2)
which is independent of $\pi$ and therefore holds for all MNEs.
The optimal $\theta_{h}$ in (2) implies that MNEs of different sizes share the same tax base elasticity. Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, independent of $\pi$. Equal tax base elasticities remove countries’ incentives for tax discrimination across MNEs in the absence of the GMT. This allows us to focus on the splitting responses to a GMT that is applied only to large firms.
The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive. Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity. However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs). [^13] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms. Applying a similar empirical strategy to US firms, Coles et al., (2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size. See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes. Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark.
1.3 Governments’ tax setting choices
Before examining the effects of the GMT, we first solve the unconstrained optimisation problem in the absence of the GMT. The non-haven country $n$ maximises a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, whose owners reside exclusively in the non-haven. The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, , 2023). For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues (e.g., Janeba and Schjelderup, , 2023; Hindriks and Nishimura, , 2025).
The government of the non-haven $n$ therefore maximises
$G_{n}$ $=\underbrace{\int_{\underline{\pi}}^{\infty}\widetilde{\pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)dF}_{\text{Private benefit}}+\lambda\underbrace{\int_{\underline{\pi}}^{\infty}t_{n}\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi dF}_{\text{Tax revenues}}$
$=(1-t_{n})\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi+\sum_{h=1}^{H}(1-t_{h})\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi-\sum_{h=1}^{H}\frac{\delta}{2}\left(\frac{t_{n}-t_{h}}{\delta}\right)^{2}\Pi$
$\hskip 18.49988pt\hskip 18.49988pt+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ is the total true profits of MNEs.
Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
$G_{h}=\lambda\int_{\underline{\pi}}^{\infty}t_{h}\theta_{h}\pi dF=\lambda t_{h}\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi.$ (4)
To determine world welfare, we take a utilitarian approach and define $G_{W}\equiv G_{n}+\sum_{h=1}^{H}G_{h}$, using (1.3) and (4).
Solving the first-order conditions of the two sets of countries yields best responses: [^14] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
$t_{n}=\frac{(\lambda-1)\left(\sum_{h=1}^{H}t_{h}+\delta\right)}{H(2\lambda-1)},\qquad t_{h}=\frac{t_{n}}{2}.$ (5)
Solving for the optimal tax rates in the unconstrained benchmark, which we call ‘Regime 0,’ gives
$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)},\qquad t_{h}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}.$ (6)
We hereafter assume $\delta\in(0,\,3H/2)$ to ensure $t_{n}^{0}\in(0,1)$.
Two comments on (6) are in order. First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$. Second, they are rising in the MVPF, $\lambda$. A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the value of tax revenues vis-à-vis the post-tax profits of MNEs. The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well. [^15] Note the difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT. Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country. By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs; hence MNEs diversify their allocation of shifted profits across all havens.
The share of shifted profits in Regime 0 is given by
$\sum_{h=1}^{H}\theta_{h}^{0}=H\cdot\left(\frac{t_{n}^{0}-t_{h}^{0}}{\delta}\right)=\frac{H}{\delta}\cdot\frac{\delta(\lambda-1)}{H(3\lambda-1)}=\frac{\lambda-1}{3\lambda-1}<\lim_{\lambda\to\infty}\sum_{h=1}^{H}\theta_{h}^{0}=\frac{1}{3}.$ (7)
Hence, when the non-haven maximises tax revenues only ($\lambda\to\infty$), the international tax differential between the non-haven and the havens is maximal, resulting in one-third of all profits of MNEs being shifted to the havens in equilibrium. For lower values of $\lambda$, the equilibrium tax differential is reduced, and the share of shifted profits is accordingly lower.
2 Theoretical results
2.1 Tax regimes
With a GMT in place, in the first stage of the game each country decides whether to commit to a single tax rate—equal to or above the GMT rate. Non-committing countries split rates in the second stage, setting different rates for large and small MNEs. Large MNEs with profits $\pi\in[\pi_{M},\infty)$ are subject to the GMT, while small MNEs with $\pi\in[\underline{\pi},\pi_{M})$ are not. Let $\phi$ denote the exogenous share of profits earned by large MNEs, i.e., the coverage rate of the GMT:
$\phi\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\int_{\underline{\pi}}^{\infty}\pi dF}\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\Pi}.$ (8)
Table 1. Possible Regimes of the Non-Cooperative Tax Game.
Non-haven Haven
single non-GMT rate single GMT rate split tax rate
single non-GMT rate Regime 0 Regime 1 Regime 2
single GMT rate Regime 3
split tax rate Regime 4
Table 1 shows possible combinations of tax policies of the non-haven and the havens. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these regimes, if the non-haven adopts a single GMT rate, the havens set a uniform tax rate at the same level or higher and therefore earn no positive tax revenues. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven for small MNEs. Alternatively, if the havens set a single non-GMT or a single GMT rate, the non-haven never wants to split its tax rate, due to the constant tax base elasticity. Hence only five possible tax regimes remain.
Regime 0 is the unconstrained tax-competition equilibrium, where the GMT rate $t_{M}$ is non-binding in all countries, as discussed earlier. As $t_{M}$ rises, the equilibrium moves through four additional regimes, characterised as follows.
Proposition 1
(Equilibrium regimes with binding GMT)
Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$. As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterised by the following four regimes, with all symmetric havens choosing the same tax schedule.
  • (i)
    Regime 1: $t_{M}\in\left[t_{M}^{0},t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\right]$.  The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
    $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
  • (ii)
    Regime 2: $t_{M}\in\left(t_{M}^{1},t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}\right]$.  The non-haven sets a single non-GMT rate, and the haven splits its tax rate and chooses the GMT rate for large MNEs, but a lower rate than the GMT for small MNEs:
    $t_{n}^{2}=\frac{2(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}\ \ \text{for}\ \pi\in[\pi_{M},\infty),\qquad t_{h}^{2}=\begin{cases}\frac{(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}\ \ &\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}\ \ &\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.$
  • (iii)
    Regime 3: $t_{M}\in\left(t_{M}^{2},t_{M}^{3}\equiv\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}\right]$.  The non-haven chooses a single GMT rate, and the haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs:
    $\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$
  • (iv)
    Regime 4: $t_{M}\in(t_{M}^{3},1]$.  Both the non-haven and the haven split their tax rates and choose the GMT rate for large MNEs, but a lower tax rate for small MNEs:
    $\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$
Proof: See Appendix 1.
In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates.
In Regime 1, when the GMT rate $t_{M}$ exceeds the havens’ unconstrained rate $t_{h}^{0}$, they must apply $t_{M}$ to large MNEs. The non-haven’s unconstrained rate remains above $t_{M}$, so it continues to set a single rate above $t_{M}$ for all MNEs. The havens then choose whether to apply $t_{M}$ also to small MNEs or use a split schedule. By committing to a single GMT rate, they can induce the non-haven to raise its uniform rate. This argument is strongest when the havens decide collectively, as assumed in our main model. [^16] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Supplemental Appendix B. In this case, the parameter range of Regime 1 is smaller, since the commitment of a single haven induces the non-haven to raise its tax rate less than the simultaneous commitment of multiple havens. The havens trade off gains from inducing a higher non-haven rate against revenue losses from not undercutting the non-haven’s rate even more for small MNEs. In Regime 1, the first effect dominates, so they commit to the GMT rate for all MNEs, leading all countries to set higher uniform tax rates than in Regime 0.
In Regime 2, with $t_{M}$ above the regime-switching rate $t_{M}^{1}$, the non-haven remains unconstrained and sets a single rate above $t_{M}$ for all MNEs. For the havens, however, the trade-off described in Regime 1 above changes. The higher $t_{M}$ reduces gains from a uniform rate, as the tax differential and thus shifted profits shrink. [^17] This is seen from Proposition 1(ii), where the tax differential $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases with $t_{M}$. At the same time, it increases losses from not competing with the non-haven for small MNEs. Consequently, the havens split their rates, applying $t_{M}$ to large MNEs and a lower rate to small MNEs.
In Regime 3, with $t_{M}$ above the regime-switching rate $t_{M}^{2}$, the non-haven can no longer set a rate above $t_{M}$ but is bound by it for large MNEs. By committing to $t_{M}$ for all MNEs, it induces the havens to raise their rates for small MNEs, analogous to the argument in Regime 1. In equilibrium, the non-haven applies $t_{M}$ to all firms, while the havens set $t_{M}/2$ for small MNEs.
Finally, in Regime 4, with $t_{M}$ above the regime-switching rate $t_{M}^{2}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set $t_{M}$ for all MNEs. [^18] This is seen from Proposition 1(iii), where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases with $t_{M}$. To limit tax base losses from small MNEs, the non-haven abandons its commitment and lowers its tax rate for them. Again, this argument is analogous to the one for the havens to end their commitment in Regime 2. In Regime 4, both countries split their rates, applying unconstrained rates to small MNEs and the GMT rate to large MNEs.
Fig.1. Equilibrium Tax Rates for Different GMT Rates.
Notes: The figure shows the equilibrium tax rates in the non-haven in panel (a) and the haven in (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes in (c). While the levels of tax rates and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.
Figure 1 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$. The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b). The equilibrium regimes are summarised in panel (c). The solid (dotted) curves are the tax rates for large MNEs (small MNEs) in the tax-competition equilibrium under the GMT; the dashed horizontal lines are the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0). Tax rates in both countries generally rise as $t_{M}$ increases, but there are distinct patterns in the different regimes. At the switch from Regime 1 to 2, the havens begin to split their tax rates, and the rate for small MNEs falls discontinuously. At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0. The havens respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
2.2 Welfare effects of the GMT
We briefly summarise how the GMT introduction affects national and global welfare. Introducing a GMT rate $t_{M}$ slightly above the havens’ unconstrained one $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the havens. At the same time, profit shifting is reduced in comparison to the unconstrained Regime 0 due to a smaller tax differential. The sum of these effects must benefit the non-haven. It also benefits the havens, as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
Once $t_{M}$ exceeds a certain threshold within Regime 1, the relationship reverses for the havens, as the loss of shifted profits dominates. Introducing any GMT rate above this threshold benefits the non-haven but harms the havens, with the net effect raising global welfare. Supplemental Appendix C provides a formal proposition and proof of these results.
We also study the effects of marginal increases in the GMT tax rate $t_{M}$ and the GMT coverage rate $\phi$. An important result is that welfare in the non-haven and the world increases with $t_{M}$ in general, and the welfare gains become even larger as $\phi$ increases. A higher $\phi$ reduces the effectiveness of the split tax system and limits tax competition for small MNEs, thereby amplifying the welfare gains from raising $t_{M}$. See Supplemental Appendices D and E for formal discussions and Appendix F for graphical illustrations.
3 Quantitative implications
We now explore the quantitative importance of our theoretical results. We modify the model of the previous section by assuming that the havens decide tax policies in a decentralised manner (see footnote 16 and Supplemental Appendix B). The decentralised decision making leads to competition between the havens and thus makes it more likely that they split their tax rates. [^19] The new regime switching rate dividing Regimes 1 and 2 is smaller ($t_{M}^{b}<t_{M}^{1}$), and thus Regime 1 is less likely to emerge, as $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$. This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens, making gains from commitment smaller.
We calibrate the unconstrained model with no GMT (Regime 0) to match international profit-shifting data, as summarised in Table 2. The GMT coverage rate is set to $\phi=0.9$, following OECD, (2020) and our calculations from Orbis. [^20] See Supplemental Appendix A for details. The number of tax havens is set to $H=40$, as in Tørsløv et al., (2023). Combining their data with Country-by-Country Reporting (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD. The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the model-predicted non-haven tax rate and shifted profit share and their empirical counterparts. For non-targeted moments, our model slightly underestimates the havens’ tax rate, $t_{h}^{0}$, and the non-haven’s revenue losses, $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$, but overall fits the data reasonably well.
Table 2. Calibration of the Model.
Parameter Source
$\phi=0.9$ Profit share of MNEs covered by GMT OECD, (2020), Orbis
$H=40$ Number of tax havens Tørsløv et al., (2023)
$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
$\lambda=2.12$ Marginal value of public funds Calibrated
$\delta=17.80$ Cost of profit shifting Calibrated
Targeted moment Model Data Source
Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al., (2023)
Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al., (2023), CbCR
Non-targeted moment Model Data Source
Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al., (2023)
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al., (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
Using this calibrated version of our model, we examine the quantitative effects of GMT reforms, as summarised in Table 3. The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields the equilibrium in Regime 2 (Table 3(a)): the havens split their tax rates and set a lower tax rate ($10.1\%$) for small MNEs as compared to large MNEs ($15.0\%$) (Proposition 1(ii)). [^21] Above the regime switching rate of $t_{M}^{b}=0.094$, Regime 2 prevails. This result aligns with observations that low-tax countries like Ireland plan to retain their base rate (12.5% in Ireland’s case) while applying top-up taxes on large MNEs under the GMT (QDMTT).
Table 3. Quantitative Effects of GMT Reforms.
(a) Gains from a 15% GMT under $\phi=0.9$ Non-haven Haven World
Tax rate after GMT introduction $20.3\%$ for all $\begin{cases}10.1\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
Welfare change in bUSD (% change) $87.2$ ($3.5\%$) 3.8-3.8 ($4.3\%$) $83.4$ ($3.2\%$)
Revenue change in bUSD (% change) $139.5$ ($20.6\%$) 3.8-3.8 ($4.3\%$) $135.7$ ($17.7\%$)
Profit change in bUSD (% change) $359.9$ $(37.5\%)$ 359.9-359.9 $(-37.5\%)$ 0 $(0\%)$
(b) Gains from a 15% GMT under $\phi=1.0$ Non-haven Haven World
Tax rate after GMT introduction $20.6\%$ for all $15\%$ for all —–
Welfare change in bUSD (% change) $95.6$ ($3.8\%$) 3.6-3.6 ($4.1\%$) $92.0$ ($3.6\%$)
Revenue change in bUSD (% change) $153.0$ ($22.6\%$) 3.6-3.6 ($4.1\%$) $149.3$ ($19.5\%$)
Profit change in bUSD (% change) $393.0$ $(41.0\%)$ 393.0-393.0 $(-41.0\%)$ 0 $(0\%)$
Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$) and panel (b), from $t_{M}=0$ to $0.15$ with full coverage $\phi=1.0$. The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven. Results are based on the calibrated parameter values described in Table 2. The welfare levels are normalised by the MVPF, $\lambda$. The ‘haven’ here refers to the sum of the $H=40$ haven countries.
The non-haven’s revenue rises due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition. The calibrated revenue gain in the non-haven (139.5 bUSD) is close to the combined gains for Europe and the US in Baraké et al., (2022, Table 1) ($67+58=125$ bEUR), and the global revenue gain (135.7 bUSD) is comparable to Hugger et al., (2024, Section 8.2) (155–192 bUSD). Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues. [^22] In the havens without MNE headquarters, revenues coincide with welfare.
Do the quantitative effects of introducing GMT change if the coverage is extended? The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) yields the non-splitting equilibrium in Regime 1 (Table 3(b)). By eliminating competition for small MNEs, the extended coverage increases revenue and welfare gains worldwide by 13.6 bUSD and 8.6 bUSD, respectively. Based on the estimates by Gaul et al., (2022), these costs would amount to about 10.5 bUSD for our sample, implying a small net revenue gain (+3.1+3.1 bUSD) but a net welfare loss (1.9-1.9 bUSD) globally. [^23] Using a survey of German MNEs (see Becker et al., , 2012 for details on their characteristics), Gaul et al., (2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR. Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in Supplemental Appendix A), the total compliance costs of newly covered MNEs are 1.6 mEUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 bUSD. Clifford et al., (2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.
4 Conclusion
This paper has analysed the effects of a global minimum tax (GMT) confined to large multinational enterprises (MNEs), leaving about 10% of the global tax base outside its scope. This partial-coverage GMT generates a sequence of tax competition equilibria between a non-haven and tax havens. In particular, introducing a low GMT rate that binds the havens induces them to commit to the single GMT rate. However, a further increase in the GMT rate leads first the havens, and then the non-haven, to split their tax rates for large and small MNEs.
The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the havens split their tax rates and undercut the GMT for small MNEs. This finding corresponds to the observation that tax havens such as Ireland and Switzerland maintain their base corporate tax rates below 15%, but top up taxes on large MNEs to the level of the GMT. In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the havens lose, resulting in a positive net gain globally.
Appendix
Appendix 1: Proof of Proposition 1
Regime 0 can be an equilibrium as long as the haven’s unconstrained rate is greater than the GMT rate, $t_{M}<t_{h}^{0}=t_{M}^{0}\equiv\delta(\lambda-1)/[H(3\lambda-1)]$. A new equilibrium regime emerges when $t_{M}\geq t_{M}^{0}$. Since haven countries are all symmetric, we are concerned with a representative haven country $h$ in what follows.
Regime 1: $t_{M}\in[t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
When the non-haven sets a single non-GMT rate and the haven sets a single GMT rate, the equilibrium tax rates and payoffs are
$t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)},\qquad t_{h}^{1}=t_{M}.$ (A.1)
$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\hskip 18.49988ptG_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
When the non-haven sets a single non-GMT rate and the haven splits its tax rate, the equilibrium tax rates are
$t_{n}^{2}=\frac{2(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]},\qquad t_{h}^{2}=\begin{cases}\frac{(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.$ (A.2)
$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\hskip 18.49988ptG_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
Given the non-haven’s single non-GMT rate, the haven prefers to set a single GMT rate, if
$G_{h}^{1}-G_{h}^{2}=$ (A.3)
$\frac{\lambda\overbrace{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]}^{\geq 0}\overbrace{\left[\delta(\lambda-1)(2\lambda-1)-Ht_{M}\left\{\lambda(3\lambda-1)-\phi(\lambda-1)^{2}\right\}\right]}^{\geq 0}}{\delta H^{2}(2\lambda-1)[\lambda(3+\phi)-(1+\phi)]^{2}}(1-\phi)\Pi\geq 0,$
where the strict equality holds at $t_{M}\in\{t_{M}^{0},t_{M}^{1}\}$. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate, and the haven sets a single GMT rate, given by (A.1).
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
Since (A.3) becomes negative for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, the haven splits its tax rate in response to a non-haven’s single non-GMT rate. The non-haven is still not bound by the GMT and continues to choose its unconstrained optimal tax rate. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate and the haven splits its tax rate (see (A.2)).
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For $t_{M}>t_{M}^{2}$, the non-haven’s single-non GMT rate, given by (A.2), is bound by the GMT and it cannot choose the unconstrained maximizing rate in response to the haven’s splitting rates. The non-haven thus chooses either a single GMT rate or a splitting tax schedule. In response, the haven must always split its tax rate; otherwise it cannot obtain positive tax revenues.
When the non-haven sets a single GMT rate and the haven splits its tax rate, the equilibrium tax rates and payoffs are
$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\hskip 18.49988ptt_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\hskip 18.49988ptG_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
When both the non-haven and the haven split their tax rates, the equilibrium tax rates are
$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\hskip 18.49988ptt_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\hskip 18.49988ptG_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
Given the haven’s splitting rates, the non-haven prefers to set a single GMT rate, if
$\begin{split}&G_{n}^{3}-G_{n}^{4}=\frac{\overbrace{\left[Ht_{M}(3\lambda-1)-2\delta(\lambda-1)\right]}^{>0}\overbrace{\left[2\delta(\lambda-1)(8\lambda-3)-Ht_{M}(3\lambda-1)(4\lambda-1)\right]}^{\geq 0}}{8\delta H(3\lambda-2)^{2}}(1-\phi)\Pi\geq 0,\end{split}$ (A.6)
which is always fulfilled in Regime 3. The strict equality holds at $t_{M}=t_{M}^{3}$ and it holds that $t_{M}>t_{M}^{2}>\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ for any $\lambda>1/3$. Therefore, the tax-competition equilibrium is that the non-haven sets a single GMT rate, and the haven splits its tax rate, as given in (A.4).
Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
Since (A.6) becomes negative for $t_{M}\in(t_{M}^{3},1]$, the non-haven splits its tax rate in response to the haven’s split tax rates. Therefore, the tax-competition equilibrium is that both the non-haven and the haven split their tax rates. ∎
Supplemental Appendix of
‘A Global Minimum Tax for Large Firms Only:
Implications for Tax Competition’
Andreas Haufler [^1] Seminar for Economic Policy, LMU Munich, Akademiestr. 1, 80799 Munich, Germany. Email: Andreas.Haufler@econ.lmu.de.     Hayato Kato [^2] Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan. Email: hayato.kato@econ.osaka-u.ac.jp
October, 2025
Contents
Appendix A MNEs Subject to the GMT
We select from the Orbis database MNEs satisfying the following conditions:
  • Global Ultimate Owner with foreign subsidiaries. The threshold ownership is 50.01%.
  • C1: MNEs report only consolidated accounts, not unconsolidated accounts.
  • Non-missing revenues (‘Operating revenue (Turnover)’) and pre-tax profits (‘P/L before tax’).
Column (a) of Table A.1 reports the pre-tax profits and the total number of MNEs in each year. Column (b) restricts the sample to MNEs subject to the GMT, i.e., those with annual revenues of at least 750 million EUR in two or more of the past four years. Using columns (a) and (b), column (c) shows the share of MNEs subject to the GMT in terms of the pre-tax profits and the number of MNEs.
Table A.1. MNEs in the Orbis Database.
(a) All MNEs (b) MNEs $\geq$ 750mEUR (c) Share == B/A
Year Pre-tax profits Number Pre-tax profits Number Pre-tax profits Number
2018 2320 8656 2066 2333 0.89 0.27
2019 2430 8416 2216 2437 0.91 0.29
2020 1985 7920 1793 2228 0.90 0.28
2021 3564 8803 3239 2627 0.91 0.30
Source: Orbis database, own calculations.
Note: Pre-tax profits are in billion EUR.
Appendix B Decentralized Decision-Making by Tax Havens
In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate. Here, we consider an alternative, decentralized decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others. In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions. First, countries can split their tax rates only if they are subject to the GMT. This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019). Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates. This allows us to focus on the problem of a single haven.
Out of the total $H$ tax havens, $H_{M}$ havens commit to a single GMT rate, while the remaining $H-H_{M}$ havens split their tax rates. When the non-haven sets a single non-GMT rate and a haven $h$ sets a single GMT rate, their equilibrium tax rates are respectively
$\begin{split}&t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\\ &t_{h}^{1}(H_{M})=t_{M},\end{split}$
Let $G_{h}^{1}(H_{M})$ denote the equilibrium payoff of the haven $h$ with commitment. We note that as in the text, setting a single non-GMT rate is the non-haven’s dominant strategy because if $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, it is never constrained by the GMT: $t_{n}(H_{M})\leq t_{n}(H)=t_{n}^{2}\leq t_{M}$.
If instead the haven $h$ splits its tax rate while the other havens and the non-haven keep their rates unchanged, the number of havens that maintain commitment falls to $H_{M}-1$. The equilibrium tax rates of the non-haven and the haven $h$ are respectively
$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (B.1)
Let $G_{h}^{2}(H_{M}-1)$ be the equilibrium payoff of the haven $h$ under splitting.
The difference between the haven $h$’s payoffs in the two cases is
$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
$\ \ =\frac{(1-\phi)\Pi[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta(H_{M})}{\delta\left[\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})\right]\left[(\lambda-1)\left\{\phi(H-H_{M}+1)+H+H_{M}-1\right\}+2H\lambda\right]^{2}},$
where
$\Theta(H_{M})\equiv AH_{M}+B,$
$A\equiv(\lambda-1)(1-\phi)\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right],$
$B\equiv H[\phi(\lambda-1)+3\lambda-1]\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right]-2t_{M}(\lambda-1)^{2}(1-\phi).$
If $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$, the haven $h$ always sets a single GMT rate, regardless of how many other havens it believes set a single GMT rate. Under our assumption of symmetric belief formation, all havens end up setting a single GMT rate, and the equilibrium coincides with Regime 1 in the text. Conversely, if $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$, all havens end up splitting their tax rates, and the equilibrium coincides with Regime 2 in the text.
Since the denominator of $\Delta G_{h}^{12}(H_{M})$ is positive and $Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0$ for $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, the sign of $\Delta G_{h}^{12}(H_{M})$ is determined by that of $\Theta(H_{M})$. Considering that $\Theta(H_{M})$ is linear in $H_{M}$, the conditions for $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$ reduce to the following two conditions:
$\Theta(1)=A\cdot 1+B=C-Dt_{M}\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{a}\equiv C/D,$
$\Theta(H)=AH+B=2(E-Ft_{M})\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{b}\equiv E/F,$
where
$C\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
$D\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
$F\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
$t_{M}^{a}-t_{M}^{0}=\frac{2\delta(\lambda-1)^{2}[\phi(\lambda-1)+3\lambda-1]}{D(3\lambda-1)}>0,$
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}[\phi(H+1)(\lambda-1)+\lambda(3H-1)-(H-1)]}{F[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0.$
We can also confirm $\Theta(H_{M})=AH_{M}+B$ increases with $H_{M}$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$, or equivalently $A>0$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$. Since $A>0$ if $t_{M}<t_{M}^{c}\equiv\delta(\lambda-1)/[H(3\lambda-1)-2(\lambda-1)]$, it suffices to show $t_{M}^{b}<t_{M}^{c}$:
$t_{M}^{c}-t_{M}^{b}=\frac{\delta(\lambda-1)[\lambda^{2}+(H-1)(2\lambda-1)-\phi(\lambda-1)^{2}]}{F[H(3\lambda-2)-(H-1)]}>0,$
which is indeed positive.
Using these results, we can thus conclude that (i) if $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$; (ii) if $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$; and (iii) if $t_{M}\in(t_{M}^{a},t_{M}^{b}]$, there exists a positive integer $H_{M}^{*}\in\{2,\dots,H\}$ such that $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,\dots,H_{M}^{*}-1\}$ and $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{H_{M}^{*},\dots,H\}$.
In cases (i) and (ii), the equilibrium is unique. In case (iii), there can be at most three equilibria. First, if each haven believes that fewer than $H_{M}^{*}$ havens set a single GMT rate, all havens end up splitting their tax rates, i.e., $H_{M}=0$. Second, if each haven believes that more than $H_{M}^{*}-1$ havens set a single GMT rate, all havens end up doing so, i.e., $H_{M}=H$. Third, if $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, then $H_{M}=H_{M}^{*}$ havens set a single GMT rate while the remaining $H-H_{M}^{*}$ havens split their tax rates, consistent with their beliefs.
The discussion is summarized as follows. If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)). If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)). If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge. If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (B.1)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (B.1)).
A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralized manner, implying that they are less likely to commit to a single GMT rate. This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
Appendix C Welfare Effects of the Introduction of GMT
The following proposition characterizes the welfare effects of introducing a binding GMT rate, compared with the unconstrained equilibrium (Regime 0).
Proposition C.1
(GMT introduction)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text, and assume the GMT coverage rate is sufficiently high, such that $\phi\in(3/4,1)$. Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
  • (i)
    a rise in welfare in the non-haven for all Regimes 1 to 4.
  • (ii)
    a rise in welfare in the (representative) haven for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
  • (iii)
    a rise in world welfare for all Regimes 1 to 4.
These results hold for any value of the marginal value of public funds (MVPF), $\lambda\in(1,\infty)$. In the limit as $\lambda\to\infty$, countries care only about tax revenues. Hence, qualitatively identical results apply to tax revenues.
Proof: In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
Comparing the welfare levels of countries in Regime 1 with those in Regime 0 gives
$G_{n}^{1}-G_{n}^{0}=\frac{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[Ht_{M}\lambda^{2}(3\lambda-1)+\delta(\lambda-1)(7\lambda^{2}-8\lambda+2)\right]}{2\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi>0,$
$G_{h}^{1}-G_{h}^{0}=\frac{\lambda\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[\delta(\lambda-1)(2\lambda-1)-Ht_{M}\lambda(3\lambda-1)\right]}{\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi\equiv\Delta$
$\Delta\ \begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\equiv t_{M}^{+}\\ <0&\text{if}\ t_{M}>t_{M}^{+}\end{cases},$
$G_{W}^{1}-G_{W}^{0}=\frac{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[\delta(\lambda-1)(11\lambda^{2}-10\lambda+2)-Ht_{M}\lambda^{2}(3\lambda-1)\right]}{2\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi>0,$
where the last inequality holds from
$t_{M}\leq t_{M}^{1}<\frac{\delta(\lambda-1)(11\lambda^{2}-10\lambda+2)}{H\lambda^{2}(3\lambda-1)}.$
In Regime 1, the welfare effect of the introduction a GMT rate is positive in the non-haven and the world. In the haven, it is positive if $t_{M}\in[t_{M}^{0},t_{M}^{+})$ and negative if $t_{M}\in(t_{M}^{+},t_{M}^{1}]$.
A marginal increase in $t_{M}$ leads to
$\frac{\partial G_{n}^{1}}{\partial t_{M}}=\frac{H\lambda^{2}t_{M}+\delta(\lambda-1)^{2}}{\delta(2\lambda-1)}\Pi>0,$
$\frac{\partial G_{h}^{1}}{\partial t_{M}}=\frac{\lambda[\delta(\lambda-1)-2H\lambda t_{M}]}{\delta(2\lambda-1)}\Pi\begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)}{2H\lambda}\equiv t_{M}^{++}\\ <0&\text{if}\ t_{M}>t_{M}^{++}\end{cases},$
$\frac{\partial G_{W}^{1}}{\partial t_{M}}=\frac{\delta(\lambda-1)(2\lambda-1)-H\lambda^{2}t_{M}}{\delta(2\lambda-1)}\Pi>0,$
where the last inequality holds because of
$t_{M}\leq t_{M}^{1}<\frac{\delta(\lambda-1)(2\lambda-1)}{2H\lambda}.$
In Regime 1, the welfare effect of a marginal increase in the GMT rate is positive in the non-haven and the world. In the haven, it is positive if $t_{M}\in[t_{M}^{0},\delta(\lambda-1)/(2H\lambda))$ and negative if $t_{M}\in(\delta(\lambda-1)/(2H\lambda),t_{M}^{1}]$.
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
For the non-haven, we see
$\frac{\partial G_{n}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{1}}=\frac{(\lambda-1)[8\lambda^{3}-5\lambda^{2}-2\lambda+1+\phi(\lambda-1)^{2}]}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi>0,$
$\frac{\partial G_{n}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)[8\lambda^{2}-5\lambda+1+\phi(\lambda-1)]}{[3\lambda-1+\phi(\lambda-1)][3\lambda-1-\phi(\lambda-1)]}\phi\Pi>0.$
From these results and the fact that $G_{n}^{2}$ is a quadratic function of $t_{M}$, it follows that $G_{n}^{2}$ increases with $t_{M}$ for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$. To prove $G_{n}^{2}-G_{n}^{0}>0$, it is sufficient to show
$G_{n}^{2}-G_{n}^{0}\big|_{t_{M}=t_{M}^{1}}=\frac{\delta(\lambda-1)^{3}[(3\lambda-1)(16\lambda^{3}-13\lambda^{2}+1)-\phi(\lambda-1)^{3}(8\lambda-3)]}{2H(3\lambda-1)^{2}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]^{2}}\phi\Pi>0,$
which holds for any $\lambda>1$ and $\phi\in(0,1)$.
For the haven, we get
$\frac{\partial G_{h}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{1}}=-\frac{2\lambda(2\lambda-1)(\lambda-1)^{2}}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi<0,$
$\frac{\partial G_{h}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=-\frac{4\lambda^{2}(\lambda-1)}{[3\lambda-1+\phi(\lambda-1)][3\lambda-1-\phi(\lambda-1)]}\phi\Pi<0.$
From these results and the fact that $G_{h}^{2}$ is a quadratic function of $t_{M}$, $G_{h}^{2}$ must be decreasing in $t_{M}$ for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$. To prove $G_{h}^{2}-G_{h}^{0}>0$, it is sufficient to show
$G_{h}^{2}-G_{h}^{0}\big|_{t_{M}=t_{M}^{1}}=-\frac{\delta\lambda(\lambda-1)^{5}[3\lambda-1+\phi(\lambda-1)]}{H(3\lambda-1)^{2}(3\lambda-1)^{2}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]^{2}}\phi\Pi<0,$
which holds for all $\lambda>1$ and $\phi\in(0,1)$.
For the world, we see
$\frac{\partial G_{W}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{1}}=\frac{(\lambda-1)[4\lambda^{3}+\lambda^{2}-4\lambda+1+\phi(\lambda-1)^{2}]}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi>0,$
$\frac{\partial G_{W}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)^{2}(4\lambda-1+\phi)}{[3\lambda-1+\phi(\lambda-1)][3\lambda-1-\phi(\lambda-1)]}\phi\Pi>0.$
Hence $G_{W}^{2}$ must be increasing in $t_{M}$ for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$. To prove $G_{W}^{2}-G_{W}^{0}>0$, we must show
$G_{W}^{2}-G_{W}^{0}\big|_{t_{M}=t_{M}^{1}}=\frac{\delta(\lambda-1)^{3}[(3\lambda-1)(14\lambda^{3}-9\lambda^{2}-2\lambda+1)-\phi(\lambda-1)^{3}(10\lambda-3)]}{2H(3\lambda-1)^{2}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]^{2}}\phi\Pi>0,$
which holds for all $\lambda>1$ and $\phi\in(0,1)$.
In Regime 2, both the introduction and a marginal increase of the GMT rate raise welfare in the non-haven and the world, but reduce it in the haven.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For the non-haven, we have
$\frac{\partial G_{n}^{3}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)[2\lambda-1+\phi(2\lambda+1)]}{2[3\lambda-1-\phi(\lambda-1)]}\Pi>0,$
$\frac{\partial G_{n}^{3}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{3}}=\frac{(\lambda-1)[\phi(8\lambda-3)-(2\lambda-1)]}{2(3\lambda-1)}\Pi>0,$
so that $G_{n}^{3}$ must increase with $t_{M}$ for $t_{M}\in(t_{M}^{2},t_{M}^{3}]$. To prove $G_{n}^{3}-G_{n}^{0}>0$, we show that the following must hold for all levels of $\lambda>1$ and $\phi\in(0,1)$:
$G_{n}^{3}-G_{n}^{0}\big|_{t_{M}=t_{M}^{2}}=\frac{\delta(\lambda-1)^{2}[(3\lambda-1)(16\lambda^{3}-13\lambda^{2}+3)-\phi(\lambda-1)^{2}(8\lambda-3)]}{2H(3\lambda-1)^{2}[3\lambda-1-\phi(\lambda-1)]^{2}}\phi\Pi>0.$
For the haven, we see
$\frac{\partial G_{h}^{3}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=\frac{\lambda(\lambda-1)}{3\lambda-1-\phi(\lambda-1)}(1-\phi)\Pi>0,$
$\frac{\partial G_{h}^{3}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{3}}=\frac{\lambda(\lambda-1)(8\lambda-3)}{(3\lambda-1)(4\lambda-1)}(1-\phi)\Pi>0,$
which implies that $G_{h}^{3}$ decreases with $t_{M}$ for $t_{M}\in(t_{M}^{2},t_{M}^{3}]$. To prove $G_{h}^{3}-G_{h}^{0}<0$, we must show that
$G_{h}^{3}-G_{h}^{0}\big|_{t_{M}=t_{M}^{3}}=\frac{\delta\lambda(\lambda-1)^{2}[8(2\lambda-1)(3\lambda-1)-\phi(8\lambda-3)^{2}]}{H(3\lambda-1)^{2}(4\lambda-1)^{2}}\Pi<0.$
This inequality holds under our assumptions of $\lambda>1$ and $\phi\in(3/4,1)$, noting that $\phi>\frac{3}{4}>\frac{8(2\lambda-1)(3\lambda-1)}{(8\lambda-3)^{2}}$.
Considering that both the non-haven and the haven increases their welfare along with a gradual increase in $t_{M}$, world welfare must also increase as $t_{M}$ rises. To prove $G_{W}^{3}-G_{W}^{0}>0$, we show that, for all $\lambda>1$ and $\phi\in(0,1)$, it holds that
$G_{W}^{3}-G_{W}^{0}\big|_{t_{M}=t_{M}^{2}}=\frac{\delta(\lambda-1)^{2}[(3\lambda-1)(14\lambda^{2}-15\lambda+3)-\phi(\lambda-1)^{2}(10\lambda-3)]}{2H(3\lambda-1)^{2}[3\lambda-1-\phi(\lambda-1)]^{2}}\phi\Pi>0.$
In Regime 3, the introduction of a GMT rate increases welfare in the non-haven and the world, while reducing it in the haven. A marginal increase in the GMT rate raises welfare in the non-haven, the haven and the world.
Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For the non-haven, we see
$G_{n}^{4}-G_{n}^{0}\big|_{t_{M}=t_{M}^{3}}=\frac{\delta(\lambda-1)^{2}(8\lambda-3)^{2}}{2H(3\lambda-1)^{2}(4\lambda-1)}\phi\Pi>0,$
for $\lambda>1$ and $\phi\in(0,1)$. Moreover, the marginal effect of the GMT is always positive: $\partial G_{n}^{4}/\partial t_{M}=(\lambda-1)\phi\Pi>0$. With these observations and the fact that $G_{n}^{4}-G_{n}^{0}$ is linear in $t_{M}$, we conclude that $G_{n}^{4}-G_{n}^{0}>0$ for $t_{M}\in(t_{M}^{3},1]$.
For the haven, we have
$G_{h}^{4}-G_{h}^{0}=-\frac{\delta\lambda(\lambda-1)^{2}}{H(3\lambda-1)^{2}}\phi\Pi>0,$
which holds for $\lambda>1$ and $\phi\in(0,1)$. The marginal effect of the GMT is zero: $\partial G_{h}^{4}/\partial t_{M}=0$ for $t_{M}\in(t_{M}^{3},1]$. Combining these results leads to $G_{W}^{4}-G_{W}^{0}>0$ and $\partial G_{W}^{4}/\partial t_{M}>0$.
In Regime 4, the introduction of a GMT rate increases welfare in the non-haven and the world, while reducing it in the haven. A marginal increase in the GMT rate raises welfare in the non-haven and the world, but has no effect on welfare in the haven. ∎
Appendix D Welfare Effects of Increasing the GMT Rate
The following proposition characterises the welfare effects of a marginal increase in the GMT rate $t_{M}$.
Proposition D.1
(Increase in the GMT rate $t_{M}$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text. A marginal increase in the GMT rate has the following effects:
  • (i)
    welfare in the non-haven increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it falls discontinuously.
  • (ii)
    welfare in the (representative) haven increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases thereafter in Regime 1 and in Regime 2. It increases in Regime 3, falls discontinuously at the regime-switching point $t_{M}^{3}$, and remains unchanged in Regime 4.
  • (iii)
    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
As is true for the introduction of the GMT (Proposition C.1 in Section C of this Supplemental Appendix), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium. An exception is the regime-switching point $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs. Small MNEs then shift more profits to the havens, hurting the non-haven.
For havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven. These opposing forces generate a non-monotonic welfare response. In Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{1}]$), the positive effect dominates initially (when $t_{M}<t_{M}^{++}$), but as the GMT rate rises further, the negative effect prevails (when $t_{M}\geq t_{M}^{++}$). In Regime 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$), the negative effect dominates throughout. In Regime 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$), the positive effect of a higher tax rate dominates again, as the non-haven commits to a uniform GMT rate in order to raise the havens’ tax rates. This regime, where commitment induces effective tax coordination, is similar to the first part of Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{++})$). It ends at the regime-switching point $t_{M}^{3}$ at which the non-haven starts splitting its tax rate. This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens. In Regime 4 ($t_{M}\in(t_{M}^{3},1]$), both the non-haven and the havens apply the same GMT rate to large MNEs, preventing them from shifting profits. Consequently, changes in the GMT rate no longer affect the havens’ welfare.
The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive. At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs. Therefore, tax competition is tightened and world welfare falls at the margin.
In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
Proof:  In the proof of Proposition C.1 in Section C, we already showed the welfare effects of a marginal increase in the GMT rate within each regime. Here we show the welfare effects at the regime switching points.
At the three regime switching GMT rates, $t_{M}^{1}$, $t_{M}^{2}$ and $t_{M}^{3}$, welfare in the non-haven, the haven, and the world changes as follows.
Change from Regime 1 to 2 at $t_{M}^{1}$
$G_{n}^{2}-G_{n}^{1}\big|_{t_{M}=t_{M}^{1}}=G_{W}^{2}-G_{W}^{1}\big|_{t_{M}=t_{M}^{1}}$
$=-\frac{\delta(\lambda-1)^{3}[\lambda(3\lambda-1)(3\lambda-2)-\phi(\lambda-1)^{2}]}{2H(\lambda-1)[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}(1-\phi)\Pi<0,$
$G_{h}^{2}-G_{h}^{1}\big|_{t_{M}=t_{M}^{1}}=0.$
Change from Regime 2 to 3 at $t_{M}^{2}$
$G_{n}^{3}-G_{n}^{2}\big|_{t_{M}=t_{M}^{2}}=G_{h}^{3}-G_{h}^{2}\big|_{t_{M}=t_{M}^{2}}=G_{W}^{3}-G_{W}^{2}\big|_{t_{M}=t_{M}^{2}}=0.$
Change from Regime 3 to 4 at $t_{M}^{3}$
$G_{n}^{4}-G_{n}^{3}\big|_{t_{M}=t_{M}^{3}}=0,$
$G_{h}^{4}-G_{h}^{3}\big|_{t_{M}=t_{M}^{3}}=G_{W}^{4}-G_{W}^{3}\big|_{t_{M}=t_{M}^{3}}=-\frac{8\delta\lambda(\lambda-1)^{2}(2\lambda-1)}{2H(\lambda-1)[H(3\lambda-1)(4\lambda-1)^{2}]}(1-\phi)\Pi<0.$
This completes the proof of Proposition D.1. ∎
Appendix E Welfare Effects of Increasing the GMT Coverage Rate
The following proposition characterizes the welfare effects of a marginal increase in the GMT coverage rate $\phi$.
Proposition E.1
(Increasing in the GMT coverage rate $\phi$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text. Except for special cases where a regime switch occurs, a marginal increase in the GMT coverage rate of $\phi\in(0,1)$ has the following effects:
  • (i)
    welfare in the non-haven is unaffected in Regime 1 and increases in Regimes 2 to 4.
  • (ii)
    welfare in the (representative) haven is unaffected in Regime 1 and decreases in Regimes 2 to 4.
  • (iii)
    world welfare is unaffected in Regime 1 and increases in Regimes 2 to 4.
In Regime 1, an increase in $\phi$ has no effect on welfare in any country, because all countries levy uniform tax rates. Changing $\phi$ does have effects, however, in the other regimes, where one or both countries split their tax rates. Since the equilibrium tax differential is always greater for small MNEs than for large MNEs, a higher $\phi$, which implies more large MNEs subject to the GMT, reduces aggregate profit shifting. This unambiguously hurts the haven and benefits the non-haven. From the global perspective, a higher GMT coverage is always desirable, as it reduces pressure on tax competition for the profits of small MNEs.
There are two exceptional cases where increasing $\phi$ causes a regime switch. The first is at $t_{M}=t_{M}^{1}+\epsilon$ with $\epsilon$ being a small positive number, where a rise in $\phi$ pushes Regime 2 back to Regime 1 by weakening the havens’ incentive to split their tax rates. This regime switch benefits the non-haven and the world, while it has no effect on the havens. The other is at $t_{M}=t_{M}^{2}+\epsilon$, where a higher $\phi$ triggers a switch from Regime 3 to Regime 2, inducing the non-haven to set a tax rate above the GMT. This latter regime switch, however, leaves welfare in all countries and the world unaffected.
In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition 3). A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1. One notable difference from the effects of a higher $t_{M}$ is that a higher $\phi$ can never harm the non-haven country and the world, even if it induces a regime switch.
In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
Proof:  The effects of an increase in the GMT coverage rate, $\phi$, on welfare in the non-haven, the haven, and the world are given as follows.
Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
We see
$\frac{\partial G_{n}^{1}}{\partial\phi}=0,\quad\frac{\partial G_{h}^{1}}{\partial\phi}=0,\quad\frac{\partial G_{W}^{1}}{\partial\phi}=0.$
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
For the non-haven, we see
$\frac{\partial G_{h}^{2}}{\partial\phi}=\frac{\Theta_{n}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{2\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
$\Theta_{n}\equiv\phi[Ht_{M}(\lambda-1)(16\lambda^{2}-13\lambda+3)-\delta(\lambda-1)^{2}]+Ht_{M}(3\lambda-1)^{2}+\delta(\lambda-1)(16\lambda^{2}-19\lambda+5).$
The sign of the derivative is determined by that of $\Theta_{n}$. From $\Theta_{n}(t_{M}=t_{M}^{1})>0$, $\Theta_{n}(t_{M}=t_{M}^{2})>0$, and the fact that $\Theta_{n}(t_{M})$ is linear in $t_{M}$, it follows that $\Theta_{n}>0$.
For the haven, we see
$\frac{\partial G_{h}^{2}}{\partial\phi}=\frac{\lambda\Theta_{h}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
$\Theta_{h}\equiv\phi[Ht_{M}(\lambda-1)(5\lambda-3)-\delta(\lambda-1)^{2}]-Ht_{M}(3\lambda-1)^{2}+\delta(\lambda-1)(5\lambda-3).$
The sign of the derivative is determined by that of $\Theta_{h}$. From $\Theta_{h}(t_{M}=t_{M}^{1})<0$, $\Theta_{h}(t_{M}=t_{M}^{2})<0$, and the linearity of $\Theta_{h}(t_{M})$ in $t_{M}$, it follows that $\Theta_{h}<0$.
For the world, we get
$\frac{\partial G_{W}^{2}}{\partial\phi}=\frac{\Theta_{W}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{2\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
$\Theta_{W}$ $\equiv\phi[Ht_{M}(\lambda-1)(2\lambda-1)(13\lambda-3)-\delta(\lambda-1)^{2}(2\lambda+1)]-Ht_{M}(3\lambda-1)^{2}(2\lambda-1)$
$+\delta(\lambda-1)(26\lambda^{2}-25\lambda+5).$
The sign of the derivative is determined by that of $\Theta_{W}$. From $\Theta_{W}(t_{M}=t_{M}^{1})>0$, $\Theta_{W}(t_{M}=t_{M}^{2})>0$, and the linearity of $\Theta_{W}(t_{M})$ in $t_{M}$, it follows that $\Theta_{W}>0$.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
We see
$\frac{\partial G_{n}^{3}}{\partial\phi}=\frac{H(4\lambda-1)t_{M}^{2}}{8\delta}\Pi>0,\qquad\frac{\partial G_{h}^{3}}{\partial\phi}=-\frac{H\lambda t_{M}^{2}}{8\delta}\Pi<0,\qquad\frac{\partial G_{W}^{3}}{\partial\phi}=\frac{H(2\lambda-1)t_{M}^{2}}{4\delta}\Pi>0,$
Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
We get
$\frac{\partial G_{n}^{4}}{\partial\phi}=\frac{(\lambda-1)[2Ht_{M}(3\lambda-1)^{2}-\delta(\lambda-1)(8\lambda-3)]}{2H(3\lambda-1)^{2}}\Pi>0,$
$\frac{\partial G_{h}^{4}}{\partial\phi}=-\frac{\delta\lambda(\lambda-1)^{2}}{H(3\lambda-1)^{2}}\Pi<0,$
$\frac{\partial G_{W}^{4}}{\partial\phi}=\frac{(\lambda-1)[2Ht_{M}(3\lambda-1)^{2}-\delta(\lambda-1)(10\lambda-3)]}{2H(3\lambda-1)^{2}}\Pi>0,$
noting that $t_{M}>t_{M}^{3}>\frac{\delta(\lambda-1)(10\lambda-3)}{2H(3\lambda-1)^{2}}>\frac{\delta(\lambda-1)(8\lambda-3)}{2H(3\lambda-1)^{2}}$.
This completes the proof of Proposition E.1. ∎
Appendix F Supplementary Figures
Figure F.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those under full GMT coverage (dashed curves). The equilibrium tax rates under full GMT are as follows.
Regime 0: $t_{M}\in\left[0,t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)}\right)$.  Each of the non-haven and the haven sets its own single non-GMT rate:
$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
Regime 1f: $t_{M}\in\left[t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)},t_{M}^{f}\equiv\frac{\delta(\lambda-1)}{H\lambda}\right]$.  The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
$t_{n}^{1f}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
Regime 2f: $t_{M}\in(t_{M}^{f},1]$.  Both the non-haven and the haven choose a single GMT rate:
$t_{n}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the haven under full GMT. In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
We note $t_{M}^{f}>t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$, given $\phi\in(0,1)$ and $\lambda>1$. That is, the equilibrium outcomes under partial GMT begins to diverge from those under full GMT once $t_{M}$ enters the range of Regime 1. This is particularly evident in panels (a) and (c).
Fig. F.1. Equilibrium Welfare and Shifted Profits.
Notes: The figure shows the equilibrium levels of welfare in the non-haven ($G_{n}$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$. The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0). While the levels of welfare, shifted profits and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.